OpenStudy (peterg2001):

help will medal A function follows the rule y = -75 - 5x. When the function's output is 25, the equation is 25 = -75 - 5x. What is the function input when the output is 25? 20 -20 10 -10

OpenStudy (peterg2001):

@shinebrightlikeadimon

OpenStudy (lcnettech):

i think its c but im not sure

OpenStudy (anonymous):

25 = -75 -5x +75 +75 100 = -5x /-5 /-5 -20 = x x = -20